OpenStudy (anonymous):

Given the point (3, 4) and the slope of 6, find y when x = 27. answers: a) 122 b)14 c)141 d)148

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

You should recognize the forumla for a slope: \[y-y_0=k(x-x_o)\]

OpenStudy (texaschic101):

y - y1 = m(x - x1) y - 4 = 6(x - 3) y - 4 = 6x - 18 y = 6x - 18 + 4 y = 6x - 14 Now to find y when x = 27, just sub in 27 for x. y = 6(27) - 14 I think you can handle it now

OpenStudy (anonymous):

ya thx

OpenStudy (texaschic101):

no problem :)